# Mitigating Downstream Model Risks via Model Provenance

### v1

@@ 3.1 Threats and Implications
The use of LAION-5B [32] in PLIP’s training poses significant risks. In December 2022, the Stanford Internet Observatory discovered that LAION-5B contained hundreds of known child sexual abuse material (CSAM) images  [^1] See Stanford Internet Observatory (SIO): identified hundreds of known images of CSAM in LAION-5B. Although PLIP developers used only a subset of 32,000 images, they failed to rigorously filter out CSAM, meaning it may have been part of the training data. Research indicates that models like PLIP, which do not generate new images, can still retain and potentially reconstruct training samples, propagating unethical and illegal content to downstream models [24]. Beyond CSAM, datasets like DukeMTMC [41], Labeled Faces in the Wild (LFW) [15], and MS-Celeb-1M [13] have encountered ethical issues, such as privacy violations and lack of informed consent [29]. Despite these problems and some data retractions, their downstream byproducts remain online. For instance, the LFW dataset, despite its flaws, saw over 3,000 downloads last month on platforms like Hugging Face.
@@ 5 Conclusion
Special thanks to Adriana Romero-Soriano, Ahrav Dutta, Carole-Jean Wu, Koustuv Sinha & Meg Risdal for their insightful comments and suggestions. This research was made possible by the scholarship from NSERC (Natural Sciences and Engineering Research Council of Canada).

### v2

@@ 3.1 Threats and Implications
The use of LAION-5B [32] in PLIP’s training poses significant risks. In December 2023, the Stanford Internet Observatory discovered that LAION-5B contained hundreds of known child sexual abuse material (CSAM) images  [^1] See Stanford Internet Observatory (SIO): identified hundreds of known images of CSAM in LAION-5B. Although PLIP developers used only a subset of 32,000 images, they failed to rigorously filter out CSAM, meaning it may have been part of the training data. Research indicates that models like PLIP, which do not generate new images, can still retain and potentially reconstruct training samples, propagating unethical and illegal content to downstream models [24]. Beyond CSAM, datasets like DukeMTMC [41], Labeled Faces in the Wild (LFW) [15], and MS-Celeb-1M [13] have encountered ethical issues, such as privacy violations and lack of informed consent [29]. Despite these problems and some data retractions, their downstream byproducts remain online. For instance, the LFW dataset, despite its flaws, saw over 3,000 downloads last month on platforms like Hugging Face.
@@ 5 Conclusion
Special thanks to Adriana Romero-Soriano, Ahrav Dutta, Carole-Jean W, & Koustuv Sinha for their insightful comments and suggestions. This research was made possible by the scholarship from NSERC (Natural Sciences and Engineering Research Council of Canada).